# Removal and Selection: Improving RGB-Infrared Object Detection via Coarse-to-Fine Fusion

### v4

@@ III-C Dynamic Feature Selection
For the “Fine Selection”, we implement the dynamic modality feature selection in the fusion process by employing a mixture of scale-aware experts to gate multi-scale features. As shown in Figure 4, we design a dedicated expert for each scale modality-specific feature. Then, we aggregate these features by using the gating network to predict a set of dynamic weights. Specifically, after obtaining different scale features $C_{\text{I-spe}}^{i}$ and $C_{\text{V-spe}}^{i}$ through the feature extraction for $I^{{}^{\prime}}$ and $V^{{}^{\prime}}$, we first make $C_{\text{I-spe}}^{i}$ and $C_{\text{V-spe}}^{i}$ go through the average pooling operations, and flattened them into one-dimension vector $X_{I}^{i}\in\mathbb{R}^{M}$ and $X_{V}^{i}\in\mathbb{R}^{M}$ to predict the weights $w_{I}^{i}$ and $w_{V}^{i}$ of the gating network $G$. It can be formulated as follows:
@@ III-D Removal and Selection Detector (RSDet)
To evaluate the effectiveness of our Coarse-to-Fine fusion strategy, we embed it into an existing object detection framework. In specific, we utilize a two-stage detector Faster R-CNN [47] as our base model and replace its backbone with our strategy to construct a new object detector called RSDet. Other modules such as Region Proposal Network (RPN) and RoI head remain unchanged.
Loss functions. To extract tyhe shared and specific feature from the images $I^{{}^{\prime}}$ and $V^{{}^{\prime}}$, we maximize the mutual information between $C_{\text{I-spe}}$ and $C_{\text{V-spe}}$ with $C_{\text{sha}}$ The mutual information can serve as the deep supervising loss function $\mathcal{L}_{\text{I-spe}}$ and $\mathcal{L}_{\text{V-spe}}$ to guide the shared-specific features learning. The definitions are as follows:
@@
where $\operatorname{MI}$ represents the operation of mutual information. We use cross-entropy ($\operatorname{CE}$) and KL-divergence ($\operatorname{KL}$) to approximate equivalent optimize the mutual information [48, 49] between different features in the latent space.
@@
For detection loss, we also use the $\mathcal{L}_{\text{rpn}}$, $\mathcal{L}_{\text{reg}}$ and $\mathcal{L}_{\text{cls}}$ same as the Faster R-CNN [47] to supervise the training of whole Removal and Selection Detector. Thus, the total loss is the sum of these individual losses:
@@ IV Experiments
TABLE I: Detection results ($\text{MR}^{\text{-2}}$, in%) under ‘All-dataset’ settings of different pedestrian distances, occlusion levels, and light conditions (Day and Night) on the KAIST dataset. The pedestrian distances consist of ‘Near’ (115 $\leq$ height), ‘Medium’ (45 $\leq$ height $<$ 115) and ‘Far’ (1 $\leq$ height $<$ 45), while occlusion levels consist of ‘None’ (never occluded), ‘Partial’ (occluded to some extent up to half) and ‘Heavy’ (mostly occluded). IoU $=0.5$ / $0.7$ is set for evaluation. The best results are highlighted in red and the second-place are highlighted in blue.
$\text{MR}^{\text{-2}}$ (IoU=0.5)
@@
RSDet (Ours) ResNet50 0.00 12.13 39.80 20.49 33.25 57.60 25.83 26.48 26.02
((a))
$\text{MR}^{\text{-2}}$ (IoU=0.7)
@@
RSDet (Ours) ResNet50 12.68 46.73 73.95 55.35 65.03 84.08 59.72 62.46 60.65
((b))
IV-A Datasets
Figure 5: The FPPI-MR curves comparisons with the state-of-the-art methods on the KAIST dataset under the ‘All-dataset’ settings [25].
IV-A1 KAIST Dataset
KAIST [25] is a public multispectral pedestrian detection dataset. Due to the problematic annotations in the original dataset, our framework is trained on 8963 image pairs with new annotations [52] and evaluated on 2252 image pairs with the improved annotations [17]. To comprehensively evaluate the performance of our method, we perform comparison experiments under ‘All-dataset’ settings [25] and six subset datasets in terms of different pedestrian distances containing “Near”, “Medium”, and “Far” and different occlusion levels containing “None”, “Partial”, and “Heavy”.
IV-A2 FLIR-aligned Dataset
FLIR is a paired visible and infrared object detection dataset including daytime and night scenes. Since the images are misaligned in the original dataset, we use the FLIR-aligned dataset [26]. It has 5142 aligned RGB-IR image pairs, of which 4129 are used for training and 1013 for testing, and contains three classes of objects: ’person’, ’car’, and ’bicycle’. Since there are very few instances of the ’dog’ category in the FLIR-aligned dataset, we clean the annotations and remove the ’dog’ category from the dataset.
IV-A3 LLVIP Dataset
LLVIP [27] is a strictly aligned visible and infrared object detection dataset for low-light vision. It is collected in low-light environments most of which were captured in very dark scenes. It contains 15488 aligned RGB-IR image pairs, of which 12025 images are used for training and 3463 images for testing.
IV-B Evaluation Metrics
IV-B1 Log-average Miss Rate ($\text{MR}^{\text{-2}}$)
For the KAIST dataset, we employ the standard KAIST evaluation [25]: Miss Rate (MR) over False Positive Per Image (also denoted as $\text{MR}^{\text{-2}}$). It calculates the average miss rate under the 9 FPPI values which are sampled uniformly in the logarithmic interval. The lower values indicate better performance.
IV-B2 Average Precision (AP)
For FLIR and LLVIP datasets, we employ the commonly used object detection metric Average Precision (AP). The positive and negative samples should be divided according to the correctness of classification and Intersection over the Union (IoU) threshold. The mean Average Precision (mAP) represents the average of AP under all categories. The $\text{mAP}_{50}$ metric represents the mean AP under IoU=0.50 and The mAP metric represents the mean AP under IoU ranges from 0.50 to 0.95 with a stride of 0.05. The higher value of the mAP metrics means better performance.
TABLE II: Camparison the performance (mAP, in%) on the FLIR and LLVIP dataset. The best results are highlighted in red and the second-place are highlighted in blue.
      FLIR       LLVIP
@@ IV-B2 Average Precision (AP)
      SSD [53]       VGG16       65.5       29.6       90.2       53.5
      RetinaNet [54]       ResNet50       66.1       31.5       94.8       55.1
      Cascade R-CNN [55]       ResNet50       71.0       34.7       95.0       56.8
      Faster R-CNN [47]       ResNet50       74.4       37.6       94.6       54.5
      DDQ-DETR [56]       ResNet50       73.9       37.1       93.9       58.6       IR
      SSD [53]       VGG16       52.2       21.8       82.6       39.8
      RetinaNet [54]       ResNet50       51.2       21.9       88.0       42.8
      Cascade R-CNN [55]       ResNet50       56.0       24.7       88.3       47.0
      Faster R-CNN [47]       ResNet50       64.9       28.9       87.0       45.1
      DDQ-DETR [56]       ResNet50       64.9       30.9       86.1       46.7       RGB
      Halfway fusion [17]       VGG16       71.5       35.8       91.4       55.1
      GAFF [57]       ResNet18       74.6       37.4       94.0       55.8
      ProbEn [58]       ResNet50       75.5       37.9       93.4       51.5
      LGADet [59]       ResNet50       74.5       -       -       -
      CSAA [60]       ResNet50       79.2       41.3       94.3       59.2
      RSDet (Ours)       ResNet50       81.1       41.4       95.8       61.3       RGB+IR
IV-C Implementation Details
All the experiments are implemented in the mmdetection [61] toolbox and conducted on a single GPU NVIDIA GeForce RTX 3090. We use ResNet-50 as the backbone. During training, the batch size is set to 4 and the SGD optimizer is employed with a momentum of 0.9 and the weight decay of $1\times 10^{-4}$. The input image resolution is $640\times 512$ for FLIR and KAIST and $1280\times 1024$ for LLVIP. All experiments are trained for 12 epochs with an initial learning rate of $1\times 10^{-2}$. We use random flipping with a probability of 0.5 to increase input diversity.

### v6

@@ III-C Dynamic Feature Selection
For “Fine”, we implement the dynamic modality feature selection in the fusion process by employing a mixture of scale-aware experts to gate multi-scale features leveraging its dynamic fusion mechanism to facilitate complementary fusion across different scales. As shown in Figure 3, we design a dedicated expert for each scale modality-specific feature. Then, we aggregate these features by using the gating network to predict a set of dynamic weights. Specifically, after obtaining different scale features $C_{\text{I-spe}}^{i}$ and $C_{\text{V-spe}}^{i}$ through the feature extraction for $I^{{}^{\prime}}$ and $V^{{}^{\prime}}$, we first make $C_{\text{I-spe}}^{i}$ and $C_{\text{V-spe}}^{i}$ go through the average pooling operations, and flattened them into one-dimension vector $X_{I}^{i}\in\mathbb{R}^{M}$ and $X_{V}^{i}\in\mathbb{R}^{M}$ to predict the weights $w_{I}^{i}$ and $w_{V}^{i}$ of the gating network $G$. It can be formulated as follows:
@@
To illustrate the effectiveness of the DFS module, we visualise the features extracted by four different experts, as shown in Figure 6. From left to right, it can be seen that different experts focus on different scales objects. Each expert selects the desired features from two modality features ($C_{I-Spe}$ and $C_{V-Spe}$) so that the fused features $C_{Spe}$ exhibit greater prominence for object detection. The visualisation results indicate that the fusion method effectively integrates complementary features from different modalities across different scales, leading to a comprehensive representation.
@@ III-D Removal and Selection Detector (RSDet)
To evaluate the effectiveness of our Coarse-to-Fine fusion strategy, we embed it into an existing object detection framework. In specific, we utilize a two-stage detector Faster R-CNN [38] as our baseline model and replace its backbone with our strategy to construct a new object detector called RSDet. Other modules such as Region Proposal Network (RPN) and R-CNN head remain unchanged.
Loss functions. To extract the shared and specific feature from the images $I^{{}^{\prime}}$ and $V^{{}^{\prime}}$ after RSR module, we maximize the mutual information [39] between $C_{\text{I-spe}}$ and $C_{\text{V-spe}}$ with $C_{\text{sha}}$ The mutual information can serve as the deep supervising loss function $\mathcal{L}_{\text{I-spe}}$ and $\mathcal{L}_{\text{V-spe}}$ to guide the shared-specific features learning. The definitions are as follows:
@@
where $\operatorname{MI}$ represents the operation of mutual information. We use cross-entropy ($\operatorname{CE}$) and KL-divergence ($\operatorname{KL}$) to approximate equivalent optimize the mutual information between different features in the latent space.
@@
As for detection loss, we use the $\mathcal{L}_{\text{rpn}}$, $\mathcal{L}_{\text{reg}}$ and $\mathcal{L}_{\text{cls}}$ same as the Faster R-CNN [38] to supervise the detection process of RSDet. The total loss is the sum of these individual losses:
@@
TABLE I: Ablation study on each module result on the FLIR, LLVIP (mAP, in%) and KAIST ($\text{MR}^{\text{-2}}$ in%) dataset, under IoU=0.7. The best results are highlighted in bold.
FLIR LLVIP KAIST (‘All’ Setting)
RSR DFS $\text{mAP}_{50}$uparrow $\text{mAP}_{75}$uparrow mAP $\uparrow$ $\text{mAP}_{50}$uparrow $\text{mAP}_{75}$uparrow mAP $\uparrow$ $\text{MR}^{\text{-2}}$-Day $\downarrow$ $\text{MR}^{\text{-2}}$-Night $\downarrow$ $\text{MR}^{\text{-2}}$-All $\downarrow$
81.2 36.2 41.2 95.0 58.6 55.5 25.30 28.87 26.48
82.3 36.0 41.9 95.0 62.0 57.1 25.27 27.89 26.12
82.4 38.7 42.8 95.6 66.0 59.5 25.11 27.99 26.23
83.9 40.1 43.8 95.8 70.4 61.3 24.18 26.49 24.79
TABLE II: Comparsion our DFS module with other feature fusion methods under a fair experiment setting on the FLIR dataset. Uniformly utilize Faster R-CNN as the detector, with ResNet-50 as the backbone. The best results are highlighted in bold.
Index Modality Methods Param. GFLOPs $\text{mAP}_{50}$uparrow $\text{mAP}_{75}$uparrow mAP $\uparrow$
(a) RGB Faster R-CNN [38] 41.13M 75.6 64.9 21.1 28.9
(b) IR 41.13M 75.6 74.4 32.5 37.6
(c) RGB+IR Two Stream Faster R-CNN [38] Two Stream 64.61M 102.5 73.1 32.0 37.1
(d) RGB+IR + CMX [40] 305.20M 229.0 80.5 33.4 39.7
(e) RGB+IR + CFT [41] 196.93M 136.8 77.5 34.6 39.2
(F) RGB+IR + DFS (Ours) 68.52M 136.0 80.9 36.0 41.3
TABLE III: comparison of different design choices for the filter Type and the $K$ value of the Topk operation in the RSR module on the FLIR dataset. The best results are highlighted in bold.
Filter Type K $\text{mAP}_{50}$uparrow $\text{mAP}_{75}$uparrow mAP $\uparrow$
Hard Filter 300 82.7 36.4 42.0
320 83.2 38.4 43.3
340 80.8 36.0 41.6
360 81.9 36.9 42.1
380 83.5 38.8 43.3
400 82.4 38.7 42.8
Avg. 82.42 37.53 42.52
Soft Filter 300 83.8 38.2 43.2
320 83.9 40.1 43.8
340 82.6 38.8 43.1
360 82.2 37.7 42.6
380 83.5 36.6 42.7
400 82.4 38.7 42.8
Avg. 82.90 38.17 42.90
@@ IV Experiments
IV-A Experimental Setup
IV-A1 Datasets
KAIST [22] is a public multispectral pedestrian detection dataset. Due to the problematic annotations in the original dataset, further research has improved the annotations of train [42] and test dataset [14]. Our method is trained on 8,963 image pairs and evaluated on 2,252 image pairs with the improved annotations. The KAIST dataset is divided into different subsets [22]: near, medium, and far (“Scale”); none, partial, and heavy (“Occlusion”); day, night, and all” (All and Reasonable). In particular, the ‘All’ setting evaluates the model performance on all objects of the KAIST test dataset, while the ‘Reasonable’ setting only consists of not/partially occluded objects and objects larger than 55 pixels. To comprehensively evaluate the performance of our method, we perform comparison experiments under ‘All’ settings.
FLIR-aligned is a paired RGB-IR object detection dataset including daytime and night scenes. Since the images are misaligned in the original dataset, we use the FLIR-aligned dataset [43]. It has 5,142 aligned RGB-IR image pairs, of which 4,129 are used for training and 1,013 for testing, and contains three classes of objects: ’person’, ’car’, and ’bicycle’. Since there are very few instances of the ’dog’ category in the FLIR-aligned dataset, we clean the annotations and remove the ’dog’ category from the dataset.
LLVIP [44] is a strictly aligned RGB-IR object detection dataset for low-light vision. It is collected in low-light environments, and most of the data are captured in very dark scenes. It contains 15,488 aligned RGB-IR image pairs, of which 12,025 images are used for training and 3463 images for testing.
IV-A2 Evaluation Metrics
Log-average Miss Rate ($\text{MR}^{\text{-2}}$): For the KAIST dataset, we employ the standard KAIST evaluation [22]: Miss Rate (MR) over False Positive Per Image (also denoted as $\text{MR}^{\text{-2}}$). It calculates the average miss rate under the 9 FPPI values which are sampled uniformly in the logarithmic interval. The lower values indicate better performance.
mean Average Precision (mAP): For FLIR and LLVIP datasets, we employ the commonly used object detection metric Average Precision (AP). The positive and negative samples should be divided according to the correctness of classification and Intersection over the Union (IoU) threshold. The $\text{mAP}_{50}$ metric represents the mean AP under IoU=0.50 and the mAP metric represents the mean AP under IoU ranges from 0.50 to 0.95 with a stride of 0.05.
Figure 7: Visualization of RSR module intermediate output on the FLIR (left) and LLVIP (right) datasets. $V$ and $I$ represent the input RGB and Infrared images respectively, $V^{{}^{\prime}}$ and $I^{{}^{\prime}}$ represent the images with irrelevant redundant spectrum removed after the RSR module. “SNR” stands for signal-to-noise ratio. The green bounding box indicates the object (foreground), while the red bounding box indicates the background (no-object class). We have provided enlarged views of these regions to facilitate observing and comparing the changes in the object and background regions before and after the RSR module.
Figure 8: Visualization of DFS module feature fusion results on the FLIR dataset. To facilitate a clearer observation, we overlaid the features onto the original RGB or Infrared image.
Figure 9: tSNE visualization of the modality-specific and shared features. ‘w/o RSR’(a) and ‘w RSR’ (b) represent without and with RSR module.
IV-A3 Implementation Details
All the experiments are implemented in the mmdetection toolbox and conduct on the NVIDIA GeForce RTX 3090. We use the Faster R-CNN as the object detector with ResNet50 as the backbone. During the training phase, the SGD optimizer is employed with a momentum of 0.9 and a weight decay of $1\times 10^{-4}$. To facilitate the design of the DFS module, we adjusted the input image resolution of various datasets to match that of the LLVIP dataset. All experiments are most trained for 12 epochs with an initial learning rate of $1\times 10^{-2}$ for the FLIR-align and KAIST datasets, and $1\times 10^{-3}$ for the LLVIP datasets. As for data augmentation, we only use random flipping with a probability of 0.5 to increase input diversity.
IV-B Ablation Study
IV-B1 Effectiveness of Each Component
To rigorously assess the efficacy of the RSR and DFS modules, we perform an ablation study on each component within RSDet, engaging in an extensive series of experiments on the FLIR, LLVIP, and KAIST dataset, as shown in Table I. Both the RSR and DFS modules individually demonstrate consistent improvements across the FLIR, LLVIP, and KAIST datasets. Moreover, the superior results can be obtained by utilizing both RSR and DFS modules, as can be seen from the bold results in Table I. These results highlight the effectiveness of the RSR and DFS modules in enhancing the model’s detection capabilities.
IV-B2 Effectiveness of DFS Module
To validate the superiority of our proposed DFS method, we replace the DFS module with the other existing RGB-IR fusion module. We have conducted the comparative experiments under a unified backbone and detector. Specifically, we use a two-stream Faster R-CNN as the baseline and apply a direct feature addition fusion method for RGB and IR features [14]. For the comparison fusion methods from other studies, we replace the feature addition operation with their own feature fusion module and performed fair experiments on the FLIR dataset. Additionally, we also compare the models in terms of parameter quantity and inference computational complexity.
The experimental results are shown in Table II. Our method increases the number of parameters by only 3.91M compared to the baseline, yet it significantly improves detection performance, with a 7.8% increase in $\text{mAP}_{50}$, a 4.0% increase in $\text{mAP}_{75}$, and a 4.2% increase in mAP. Besides, our DFS module exhibits superior comprehensive performance while maintaining a lower number of parameters and computational complexity. Our method outperforms the second-best fusion method, even with 236.68M fewer parameters and 93G fewer FLOPS, showing a 0.4% lead in $\text{mAP}_{50}$, a 2.6% advantage in $\text{mAP}_{75}$, and a 1.6% improvement in mAP, achieving the best performance among different RGB-IR feature fusion methods.
@@ IV-B2 Effectiveness of DFS Module
TABLE IV: Detection results ($\text{MR}^{\text{-2}}$, in%) under ‘All’ settings of different pedestrian distances, occlusion levels, and light conditions (Day and Night) on the KAIST dataset. The pedestrian distances consist of ‘Near’ (115 $\leq$ height), ‘Medium’ (45 $\leq$ height $<$ 115) and ‘Far’ (1 $\leq$ height $<$ 45), while occlusion levels consist of ‘None’ (never occluded), ‘Partial’ (occluded to some extent up to half) and ‘Heavy’ (mostly occluded). IoU $=0.5$ and $0.7$ is set for evaluation. The best results are highlighted in red and the second-place are highlighted in blue. Noted: We conduct the experiments under the‘All’ setting, which evaluates the model performance on all objects of the KAIST test dataset, rather than the ‘Reasonable’ setting, which only consists of none/partially occluded objects, and objects larger than 55 pixels.
(a) IoU=0.5
Scale. Occlusion. All.
@@
RSDet (Ours) ResNet50 0.00 10.69 37.68 18.97 33.27 55.59 24.18 26.49 24.79
(b) IoU=0.7
Scale. Occlusion. All.
@@
RSDet (Ours) ResNet50 11.81 46.03 71.62 54.41 65.79 82.24 59.17 62.54 60.04
IV-B3 Comparative Analysis of Filter Designs within the RSR Module
We further explore the impact of different Filter Type designs and different values of $K$ in the RSR module. Specifically, if the positions of $M_{m_{I}}$ and $M_{m_{V}}$ not corresponding to the $\operatorname{top}K$ values are in the range of $\left[0,1\right]$, the Filter is the soft filter, And if are directly set to 0 is the hard Filter. We have conducted the experiments by adjusting the value of $K$ from 300 to 400 across different filter types on the FLIR datasets. Note that 400 is the total patch number of image, so, $K=400$ represents nothing removed. As illustrated in Table III, our detector achieves the highest detection performance with the Soft Filter when $K=320$. In addition to considering the optimal performance, we also evaluate the stability of each filter type. To this end, we calculate the average (Avg.) mAP value across different $K$ for each filter type. From Table III, on the FLIR-aligned dataset, it is evident that the Soft Filter consistently outperforms the others across all metrics, with the best result also achieved using the Soft Filter. Therefore, we select the Soft Filter and $K=320$ as the filter hyper-parameters.
TABLE V: Camparison the performance (mAP, in%) on the FLIR dataset. The best results are highlighted in red and the second-place are highlighted in blue.
FLIR
@@ IV-B3 Comparative Analysis of Filter Designs within the RSR Module
SSD [47] VGG16 65.5 29.6
RetinaNet [48] ResNet50 66.1 31.5
Cascade R-CNN [49] ResNet50 71.0 34.7
Faster R-CNN [38] ResNet50 74.4 37.6
DDQ-DETR [50] ResNet50 73.9 37.1 IR
SSD [47] VGG16 52.2 21.8
RetinaNet [48] ResNet50 51.2 21.9
Cascade R-CNN [49] ResNet50 56.0 24.7
Faster R-CNN [38] ResNet50 64.9 28.9
DDQ-DETR [50] ResNet50 64.9 30.9 RGB
Halfway fusion [14] VGG16 71.5 35.8
CFR_3 [43] VGG16 72.4 -
GAFF [51] VGG16 72.7 37.3
GAFF [51] ResNet18 74.6 37.4
CAPTM_3 [52] ResNet50 73.2 -
CMPD [10] ResNet50 69.4 -
LGADet [53] ResNet50 74.5 -
ProbEn [54] ResNet50 75.5 37.9
TINet [55] ResNet50 76.1 36.5
ICAFusion [56] ResNet50 72.0
ICAFusion [56] CSPDarkNet53 79.2 41.4
YOLOFusion [57] CSPDarkNet53 76.6 39.8
MFPT[58] ResNet50 80.0 -
CSAA [59] ResNet50 79.2 41.3
RSDet (Ours) ResNet50 83.9 43.8 RGB+IR
IV-C Visualization of Intermediate Results
IV-C1 Visualizations of the intermediate results in the RSR module
To demonstrate the effectiveness of the RSR module, we visualize the intermediate results, as illustrated in Figure 7. It is evident that the object regions remain largely unchanged after processing through the RSR module, while the removed information is predominantly concentrated in the background areas. This indicates that the RSR module adaptively filters out redundant background noise that is irrelevant to the detection task. For a more objective comparison, we calculate the signal-to-noise ratio (SNR) to quantify the differences. The results reveal a modest increase in SNR after the original image passes through the RSR module, further substantiating the efficacy of the RSR module.
IV-C2 Visualizations of the Shared, Specific and Fused Features
We provide visualizations of the shared and specific features, denoted as $C_{\text{sha}}$, $C_{\text{I-spe}}$, and $C_{\text{V-spe}}$, as depicted in Figure 8. In these visualizations, red boxes are used to highlight the objects of interest. By examining the features before and after the fusion process, we can clearly see the impact of our method. Specifically, the DFS module enhances the visibility of non-salient objects within the shared feature space, transforming them into salient features that are more distinguishable. This indicates that the fusion method effectively integrates information from different modalities, leading to a more comprehensive representation.
IV-C3 tSNE Visualizations of the RSR module
We also conduct tSNE visualizations of shared  ($C_{\text{sha}}$) and specific features ($C_{\text{I-spe}}$ and $C_{\text{V-spe}}$) on the FLIR dataset. As shown in Figure 9, it can be observed that without the RSR module, there is still a considerable mix of feature points in the shared and specific features, leading to difficulty in selecting the desired specific features in DFS. After adding the RSR module, we have noted a significant decrease in the number of mixed features, which verifies that removing redundant spectra is beneficial to feature disentanglement and, thus more effective in obtaining fused features in the DFS module.
TABLE VI: Camparison the performance (mAP, in%) on the LLVIP dataset. The best results are highlighted in red and the second-place are highlighted in blue.
LLVIP
[…]
SSD [47] VGG16 90.2 53.5
RetinaNet [48] ResNet50 94.8 55.1
Cascade R-CNN [49] ResNet50 95.0 56.8
Faster R-CNN [38] ResNet50 94.6 54.5
DDQ-DETR [50] ResNet50 93.9 58.6 IR
SSD [47] VGG16 82.6 39.8
RetinaNet [48] ResNet50 88.0 42.8
Cascade R-CNN [49] ResNet50 88.3 47.0
Faster R-CNN [38] ResNet50 87.0 45.1
DDQ-DETR [50] ResNet50 86.1 46.7 RGB
Halfway fusion [14] VGG16 91.4 55.1
GAFF [51] ResNet18 94.0 55.8
ProbEn [54] ResNet50 93.4 51.5
CSAA [59] ResNet50 94.3 59.2
RSDet (Ours) ResNet50 95.8 61.3 RGB+IR